# Video Prediction Transformers without Recurrence or Convolution

### v3

@@ 4.2 Real-world Human Motion Prediction
| Method | Paras(M) | Flops(G) | FPS | MSE | MAE | SSIM | PSNR | LPIPS |
| ConvLSTM | 15.5 | 347.0 | 52 | 125.5 | 1566.7 | 0.9813 | 33.40 | 0.03557 |
| PredNet | 12.5 | 13.7 | 176 | 261.9 | 1625.3 | 0.9786 | 31.76 | 0.03264 |
| PredRNN | 24.6 | 704.0 | 25 | 113.2 | 1458.3 | 0.9831 | 33.94 | 0.03245 |
| PredRNN++ | 39.3 | 1033.0 | 18 | 110.0 | 1452.2 | 0.9832 | 34.02 | 0.03196 |
| MIM | 47.6 | 1051.0 | 17 | 112.1 | 1467.1 | 0.9829 | 33.97 | 0.03338 |
| E3D-LSTM | 60.9 | 542.0 | 7 | 143.3 | 1442.5 | 0.9803 | 32.52 | 0.04133 |
| PhyDNet | 4.2 | 19.1 | 57 | 125.7 | 1614.7 | 0.9804 | 33.05 | 0.03709 |
| MAU | 20.2 | 105.0 | 6 | 127.3 | 1577.0 | 0.9812 | 33.33 | 0.03561 |
| PredRNNv2 | 24.6 | 708.0 | 24 | 114.9 | 1484.7 | 0.9827 | 33.84 | 0.03334 |
| SimVP | 41.2 | 197.0 | 26 | 115.8 | 1511.5 | 0.9822 | 33.73 | 0.03467 |
| TAU | 37.6 | 182.0 | 26 | 113.3 | 1390.7 | 0.9839 | 34.03 | 0.02783 |
| OpenSTL_ViT | 28.3 | 239.0 | 17 | 136.3 | 1603.5 | 0.9796 | 33.10 | 0.03729 |
| OpenSTL_Swin | 38.8 | 188.0 | 28 | 133.2 | 1599.7 | 0.9799 | 33.16 | 0.03766 |
| PredFormer | ||||||||
| Full Attention | 12.7 | 155.0 | 16 | 113.9 | 1412.4 | 0.9833 | 33.98 | 0.03279 |
| Fac-S-T | 12.7 | 65.0 | 76 | 153.4 | 1630.7 | 0.9784 | 32.30 | 0.04676 |
| Fac-T-S | 12.7 | 65.0 | 75 | 118.4 | 1504.7 | 0.9820 | 33.67 | 0.03284 |
| Binary-TS | 12.7 | 65.0 | 75 | 111.2 | 1380.4 | 0.9838 | 34.13 | 0.03008 |
| Binary-ST | 12.7 | 65.0 | 78 | 112.7 | 1386.3 | 0.9836 | 34.07 | 0.03017 |
| Triplet-TST* | 12.7 | 60.8 | 88 | 112.4 | 1406.2 | 0.9834 | 34.05 | 0.02748 |
| Triplet-STS* | 12.7 | 69.3 | 64 | 111.8 | 1410.3 | 0.9834 | 34.07 | 0.02933 |
| Quadruplet-TSST | 12.7 | 65.0 | 72 | 110.9 | 1380.3 | 0.9839 | 34.14 | 0.03069 |
| Quadruplet-STTS | 12.7 | 65.0 | 74 | 113.4 | 1405.7 | 0.9835 | 34.04 | 0.02918 |
@@
Human3.6m. The Human3.6M dataset [15] comprises 3.6 million unique human poses with their corresponding images, serving as a benchmark for motion prediction tasks. Human motion prediction is particularly challenging due to high resolution input and complex human movement dynamics. Following OpenSTL [28], we downsample the dataset from 1000×1000×3 to 256×256×3. We use four observations to predict the next four frames.
Compared to SimVP, PredFormer Quadruplet-TSST reduces MSE from 115.8 to 110.9, significantly improving human motion prediction. At the same time, PredFormer requires only 12.7M parameters, much fewer than SimVP’s 41.2M. Furthermore, its computational cost is only 65G FLOPs, less than one third of SimVP’s 197G, while maintaining a higher inference speed.
@@ 4.3 Traffic Flow Prediction
| Method | Paras(M) | Flops(G) | FPS | MSE | MAE | SSIM |
| ConvLSTM | 15.0 | 20.7 | 815 | 0.485 | 17.7 | 0.978 |
| PredRNN | 23.7 | 42.4 | 416 | 0.464 | 16.9 | 0.977 |
| PredRNN++ | 38.4 | 63.0 | 301 | 0.448 | 16.9 | 0.971 |
| MIM | 37.9 | 64.1 | 275 | 0.429 | 16.6 | 0.971 |
| E3D-LSTM | 51.0 | 98.2 | 60 | 0.432 | 16.9 | 0.979 |
| PhyDNet | 3.1 | 5.6 | 982 | 0.362 | 15.5 | 0.983 |
| PredRNNv2 | 23.7 | 42.6 | 378 | 0.383 | 15.5 | 0.983 |
| SwinLSTM | 2.9 | 1.3 | 1425 | 0.303 | 15.0 | 0.984 |
| SimVP | 13.8 | 3.6 | 533 | 0.414 | 16.2 | 0.982 |
| TAU | 9.6 | 2.5 | 1268 | 0.344 | 15.6 | 0.983 |
| OpenSTL_ViT | 9.7 | 2.8 | 1301 | 0.317 | 15.2 | 0.984 |
| OpenSTL_Swin | 9.7 | 2.6 | 1506 | 0.313 | 15.1 | 0.985 |
| PredFormer | ||||||
| Full Attention | 8.4 | 2.4 | 2438 | 0.316 | 14.6 | 0.985 |
| Fac-S-T | 8.4 | 2.2 | 3262 | 0.320 | 15.2 | 0.984 |
| Fac-T-S | 8.4 | 2.2 | 3224 | 0.283 | 14.4 | 0.985 |
| Binary-TS | 8.4 | 2.2 | 3192 | 0.286 | 14.6 | 0.985 |
| Binary-ST | 8.4 | 2.2 | 3172 | 0.277 | 14.3 | 0.986 |
| Triplet-TST | 6.3 | 1.6 | 4348 | 0.293 | 14.7 | 0.985 |
| Triplet-STS | 6.3 | 1.6 | 4249 | 0.277 | 14.3 | 0.986 |
| Quadruplet-TSST | 8.4 | 2.2 | 3230 | 0.284 | 14.4 | 0.986 |
| Quadruplet-STTS | 8.4 | 2.2 | 3259 | 0.293 | 14.6 | 0.985 |
TaxiBJ. TaxiBJ [42] includes GPS data from taxis and meteorological data in Beijing. Each data frame is visualized as a heatmap, where the third dimension encapsulates the inflow and outflow of traffic within a designated area. Following previous work [42], we allocate the final four weeks’ data for testing, utilizing the preceding data for training. Our prediction model uses four sequential observations to forecast the subsequent four frames.
Compared to SimVP, PredFormer significantly improves the accuracy of the prediction, reducing the MSE from 0.414 to 0.277 (33%) while using fewer parameters (8.4M vs 13.8M) and a lower computational cost (2.2G FLOPs vs. 3.6G FLOPs). Despite this efficiency, PredFormer also dramatically increases inference speed, with FPS rising from 533 in SimVP to 4249.
Furthermore, OpenSTL-ViT and OpenSTL-Swin, which adopt ViT-based architectures as temporal translators, achieve MSEs of 0.317 and 0.313, both worse than PredFormer’s best results. This suggests that using CNNs for the encoder and decoder provides strong inductive bias but inherently limits the model’s performance ceiling. Among the variants of PredFormer, Binary-ST and Triplet-STS achieve the best MSE of 0.277.
@@ 4.4 Weather Forecasting
| Method | Paras(M) | Flops(G) | FPS | MSE | MAE | SSIM |
| ConvLSTM | 14.9 | 136.0 | 46 | 1.521 | 0.7949 | 1.233 |
| PredRNN | 23.6 | 278.0 | 22 | 1.331 | 0.7246 | 1.154 |
| PredRNN++ | 38.3 | 413 | 15 | 1.634 | 0.7883 | 1.278 |
| MIM | 37.8 | 109.0 | 126 | 1.784 | 0.8716 | 1.336 |
| PhyDNet | 3.1 | 36.8 | 177 | 285.9 | 8.7370 | 16.91 |
| MAU | 5.5 | 39.6 | 237 | 1.251 | 0.7036 | 1.119 |
| PredRNNv2 | 23.6 | 279.0 | 22 | 1.545 | 0.7986 | 1.243 |
| SimVP | 14.8 | 8.0 | 196 | 1.238 | 0.7037 | 1.113 |
| TAU | 12.2 | 6.7 | 229 | 1.162 | 0.6707 | 1.078 |
| OpenSTL_ViT | 12.4 | 8.0 | 432 | 1.146 | 0.6712 | 1.070 |
| OpenSTL_Swin | 12.4 | 6.9 | 581 | 1.143 | 0.6735 | 1.069 |
| PredFormer | ||||||
| Full Attention | 5.3 | 17.8 | 177 | 1.126 | 0.6540 | 1.061 |
| Fac-S-T | 5.3 | 8.5 | 888 | 1.783 | 0.8688 | 1.335 |
| Fac-T-S | 5.3 | 8.5 | 860 | 1.100 | 0.6469 | 1.049 |
| Binary-TS | 5.3 | 8.6 | 837 | 1.115 | 0.6508 | 1.056 |
| Binary-ST | 5.3 | 8.6 | 847 | 1.140 | 0.6571 | 1.068 |
| Triplet-TST | 4.0 | 6.3 | 1064 | 1.108 | 0.6492 | 1.053 |
| Triplet-STS | 4.0 | 6.5 | 1001 | 1.149 | 0.6658 | 1.072 |
| Quadruplet-TSST | 5.3 | 8.6 | 802 | 1.116 | 0.6510 | 1.057 |
| Quadruplet-STTS | 5.3 | 8.6 | 858 | 1.118 | 0.6507 | 1.057 |
WeatherBench. Climate prediction is a critical challenge in spatiotemporal predictive learning. The WeatherBench [23] dataset provides a comprehensive global weather forecasting resource, covering various climatic factors. In our experiments, we utilize WeatherBench-S, a single-variable setup where each climatic factor is trained independently. We focus on temperature prediction at a 5.625∘ resolution ( grid points). The model is trained on data spanning 2010-2015, validated on data from 2016, and tested on data from 2017-2018, all with a one-hour temporal interval. We input the first 12 frames and predict the subsequent 12 frames in this setting.
@@ 4.5 Ablation Study and Discussion
We conduct ablation studies on our PredFormer model design and summarize the results in Tab 7 and Tab 8. We choose the best Triplet-STS model on TaxiBJ, and the best Fac-T-S model on WeatherBench as baselines.
@@
Model Regularization. Pure transformer architectures like ViT generally require large datasets for effective training, and overfitting can become challenging when applied to smaller datasets. In our experiments, overfitting is noticeable on WeatherBench and TaxiBJ. We experiment with different regularization techniques in Tab 8 and find that both dropout(DP) and stochastic depth (SD) individually improve performance compared to no regularization. However, the combination of the two provides the best results. Unlike conventional ViT practices, which use a linearly scaled drop path rate across different depths, a uniform drop path rate performs significantly better for our tasks. We adopt the exact regularization setting for all nine variants.
@@
Discussion for PredFormer Variants. Despite our in-depth analysis of the spatiotemporal decomposition, the optimal model is not definite due to the different spatiotemporal dependent properties of the datasets. We recommend starting with the Quadruplet-TSST model for diverse video prediction tasks, which consistently performs well across datasets and configurations. PredFormer uses fixed hyperparameters for spatial and temporal GTBs, leveraging the scalability of the Transformer architecture. By simply adjusting the number of PredFormer layers, optimal results can be achieved with minimal tuning.
@@ 5 Conclusion
In this paper, we introduce PredFormer, a pure transformer-based framework designed for video prediction. Our in-depth analysis extends the understanding of spatial-temporal transformer factorization, moving beyond existing video ViT frameworks. Through rigorous experiments across diverse benchmarks, PredFormer shows unparalleled performance and efficiency, surpassing previous models by a large margin. Our results show that: (1) Interleaved spatiotemporal transformer architectures establish new benchmarks, excelling across multiple datasets. (2) Factorized temporal-first encoders significantly outperform both full spatial-temporal attention encoders and Factorized spatial-first configurations. (3) Implementing dropout and uniform stochastic depth concurrently leads to superior performance enhancements on overfitting datasets. (4) Absolute position encoding consistently outperforms learnable alternatives across all benchmarks. In summary, PredFormer not only establishes a strong baseline for real-world applications but also provides a new paradigm for video prediction with recurrent-free and convolution-free design.
Supplementary Material
6 Problem Definition

### v4

@@ 4.2 Real-world Human Motion Prediction
| Method | Paras(M) | Flops(G) | FPS | Memory(MB) | MSE | MAE | SSIM | PSNR | LPIPS |
| ConvLSTM | 15.5 | 347.0 | 52 | 142.7 | 125.5 | 1566.7 | 0.9813 | 33.40 | 0.03557 |
| PredNet | 12.5 | 13.7 | 176 | 120.8 | 261.9 | 1625.3 | 0.9786 | 31.76 | 0.03264 |
| PredRNN | 24.6 | 704.0 | 25 | 327.2 | 113.2 | 1458.3 | 0.9831 | 33.94 | 0.03245 |
| PredRNN++ | 39.3 | 1033.0 | 18 | 402.3 | 110.0 | 1452.2 | 0.9832 | 34.02 | 0.03196 |
| MIM | 47.6 | 1051.0 | 17 | 434.8 | 112.1 | 1467.1 | 0.9829 | 33.97 | 0.03338 |
| E3D-LSTM | 60.9 | 542.0 | 7 | 548.9 | 143.3 | 1442.5 | 0.9803 | 32.52 | 0.04133 |
| PhyDNet | 4.2 | 19.1 | 57 | 67.9 | 125.7 | 1614.7 | 0.9804 | 33.05 | 0.03709 |
| MAU | 20.2 | 105.0 | 6 | 371.2 | 127.3 | 1577.0 | 0.9812 | 33.33 | 0.03561 |
| PredRNNv2 | 24.6 | 708.0 | 24 | 350.5 | 114.9 | 1484.7 | 0.9827 | 33.84 | 0.03334 |
| SimVP | 41.2 | 197.0 | 26 | 556.4 | 115.8 | 1511.5 | 0.9822 | 33.73 | 0.03467 |
| TAU | 37.6 | 182.0 | 26 | 551.8 | 113.3 | 1390.7 | 0.9839 | 34.03 | 0.02783 |
| OpenSTL_ViT | 11.0 | 142.2 | 35 | 1170.0 | 136.3 | 1603.5 | 0.9796 | 33.10 | 0.03729 |
| OpenSTL_Swin | 38.8 | 188.0 | 28 | 562.3 | 133.2 | 1599.7 | 0.9799 | 33.16 | 0.03766 |
| PredFormer | |||||||||
| Full Attention | 12.7 | 155.0 | 16 | 1120.7 | 113.9 | 1412.4 | 0.9833 | 33.98 | 0.03279 |
| Fac-S-T | 12.7 | 65.0 | 76 | 352.7 | 153.4 | 1630.7 | 0.9784 | 32.30 | 0.04676 |
| Fac-T-S | 12.7 | 65.0 | 75 | 356.7 | 118.4 | 1504.7 | 0.9820 | 33.67 | 0.03284 |
| Binary-TS | 12.7 | 65.0 | 75 | 352.7 | 111.2 | 1380.4 | 0.9838 | 34.13 | 0.03008 |
| Binary-ST | 12.7 | 65.0 | 78 | 348.7 | 112.7 | 1386.3 | 0.9836 | 34.07 | 0.03017 |
| Triplet-TST* | 12.7 | 60.8 | 88 | 352.7 | 112.4 | 1406.2 | 0.9834 | 34.05 | 0.02748 |
| Triplet-STS* | 12.7 | 69.3 | 64 | 356.7 | 111.8 | 1410.3 | 0.9834 | 34.07 | 0.02933 |
| Quadruplet-TSST | 12.7 | 65.0 | 72 | 352.7 | 110.9 | 1380.3 | 0.9839 | 34.14 | 0.03069 |
| Quadruplet-STTS | 12.7 | 65.0 | 74 | 356.7 | 113.4 | 1405.7 | 0.9835 | 34.04 | 0.02918 |
@@
Human3.6m. The Human3.6M dataset (Ionescu et al., 2014) comprises 3.6 million unique human poses with their corresponding images, serving as a benchmark for motion prediction tasks. Human motion prediction is particularly challenging due to high resolution input and complex human movement dynamics. Following OpenSTL (Tan et al., 2023b), we downsample the dataset from 1000×1000×3 to 256×256×3. We use four observations to predict the next four frames.
PredFormer achieves SOTA on Human3.6M. Compared to SimVP, PredFormer Quadruplet-TSST reduces MSE from 115.8 to 110.9, significantly improving human motion prediction. At the same time, PredFormer requires only 12.7M parameters, much fewer than SimVP’s 41.2M. Furthermore, its computational cost is only 65G FLOPs, less than one third of SimVP’s 197G, while maintaining a higher inference speed.
@@ 4.3 Traffic Flow Prediction
| Method | Paras(M) | Flops(G) | FPS | Memory(MB) | MSE | MAE | SSIM |
| ConvLSTM | 15.0 | 20.7 | 815 | 67.0 | 0.485 | 17.7 | 0.978 |
| PredRNN | 23.7 | 42.4 | 416 | 105.5 | 0.464 | 16.9 | 0.977 |
| PredRNN++ | 38.4 | 63.0 | 301 | 162.8 | 0.448 | 16.9 | 0.971 |
| MIM | 37.9 | 64.1 | 275 | 158.4 | 0.429 | 16.6 | 0.971 |
| E3D-LSTM | 51.0 | 98.2 | 60 | 240.5 | 0.432 | 16.9 | 0.979 |
| PhyDNet | 3.1 | 5.6 | 982 | 149.3 | 0.362 | 15.5 | 0.983 |
| PredRNNv2 | 23.7 | 42.6 | 378 | 106.8 | 0.383 | 15.5 | 0.983 |
| SwinLSTM | 2.9 | 1.3 | 1425 | 22.0 | 0.303 | 15.0 | 0.984 |
| SimVP | 13.8 | 3.6 | 533 | 183.9 | 0.414 | 16.2 | 0.982 |
| TAU | 9.6 | 2.5 | 1268 | 175.0 | 0.344 | 15.6 | 0.983 |
| OpenSTL_ViT | 9.7 | 2.8 | 1301 | 174.8 | 0.317 | 15.2 | 0.984 |
| OpenSTL_Swin | 9.7 | 2.6 | 1506 | 274.3 | 0.313 | 15.1 | 0.985 |
| PredFormer | |||||||
| Full Attention | 8.4 | 2.4 | 2438 | 42.3 | 0.316 | 14.6 | 0.985 |
| Fac-S-T | 8.4 | 2.2 | 3262 | 42.4 | 0.320 | 15.2 | 0.984 |
| Fac-T-S | 8.4 | 2.2 | 3224 | 42.4 | 0.283 | 14.4 | 0.985 |
| Binary-TS | 8.4 | 2.2 | 3192 | 42.4 | 0.286 | 14.6 | 0.985 |
| Binary-ST | 8.4 | 2.2 | 3172 | 42.4 | 0.277 | 14.3 | 0.986 |
| Triplet-TST | 6.3 | 1.6 | 4348 | 34.4 | 0.293 | 14.7 | 0.985 |
| Triplet-STS | 6.3 | 1.6 | 4249 | 34.4 | 0.277 | 14.3 | 0.986 |
| Quadruplet-TSST | 8.4 | 2.2 | 3230 | 42.4 | 0.284 | 14.4 | 0.986 |
| Quadruplet-STTS | 8.4 | 2.2 | 3259 | 42.4 | 0.293 | 14.6 | 0.985 |
TaxiBJ. TaxiBJ (Zhang et al., 2017a) includes GPS data from taxis and meteorological data in Beijing. Each data frame is visualized as a heatmap, where the third dimension encapsulates the inflow and outflow of traffic within a designated area. Following previous work (Zhang et al., 2017a), we allocate the final four weeks’ data for testing, utilizing the preceding data for training. Our prediction model uses four sequential observations to forecast the subsequent four frames.
PredFormer achieves SOTA on TaxiBJ. Compared to SimVP, PredFormer significantly improves the accuracy of the prediction, reducing the MSE from 0.414 to 0.277 (33%) while using fewer parameters (8.4M vs 13.8M) and a lower computational cost (2.2G FLOPs vs. 3.6G FLOPs). Despite this efficiency, PredFormer also dramatically increases inference speed, with FPS rising from 533 in SimVP to 4249.
Furthermore, OpenSTL-ViT and OpenSTL-Swin, which adopt ViT-based architectures as temporal translators, achieve MSEs of 0.317 and 0.313, both worse than PredFormer’s best results. This suggests that using CNNs for the encoder and decoder provides a strong inductive bias but inherently limits the model’s performance. Among the variants of PredFormer, Binary-ST and Triplet-STS achieve the best MSE of 0.277.
@@ 4.4 Weather Forecasting
| Method | Paras(M) | Flops(G) | FPS | Memory(MB) | MSE | MAE | SSIM |
| ConvLSTM | 14.9 | 136.0 | 46 | 69.8 | 1.521 | 0.7949 | 1.233 |
| PredRNN | 23.6 | 278.0 | 22 | 108.2 | 1.331 | 0.7246 | 1.154 |
| PredRNN++ | 38.3 | 413.0 | 15 | 165.5 | 1.634 | 0.7883 | 1.278 |
| MIM | 37.8 | 109.0 | 126 | 275.9 | 1.784 | 0.8716 | 1.336 |
| PhyDNet | 3.1 | 36.8 | 177 | 150.6 | 285.9 | 8.7370 | 16.91 |
| MAU | 5.5 | 39.6 | 237 | 56.2 | 1.251 | 0.7036 | 1.119 |
| PredRNNv2 | 23.6 | 279.0 | 22 | 110.8 | 1.545 | 0.7986 | 1.243 |
| SimVP | 14.8 | 8.0 | 196 | 194.7 | 1.238 | 0.7037 | 1.113 |
| TAU | 12.2 | 6.7 | 229 | 195.6 | 1.162 | 0.6707 | 1.078 |
| OpenSTL_ViT | 12.4 | 8.0 | 432 | 194.9 | 1.146 | 0.6712 | 1.070 |
| OpenSTL_Swin | 12.4 | 6.9 | 581 | 195.1 | 1.143 | 0.6735 | 1.069 |
| PredFormer | |||||||
| Full Attention | 5.3 | 17.8 | 177 | 185.4 | 1.126 | 0.6540 | 1.061 |
| Fac-S-T | 5.3 | 8.5 | 888 | 53.9 | 1.783 | 0.8688 | 1.335 |
| Fac-T-S | 5.3 | 8.5 | 860 | 54.9 | 1.100 | 0.6469 | 1.049 |
| Binary-TS | 5.3 | 8.6 | 837 | 53.9 | 1.115 | 0.6508 | 1.056 |
| Binary-ST | 5.3 | 8.6 | 847 | 51.9 | 1.140 | 0.6571 | 1.068 |
| Triplet-TST | 4.0 | 6.3 | 1064 | 48.9 | 1.108 | 0.6492 | 1.053 |
| Triplet-STS | 4.0 | 6.5 | 1001 | 49.9 | 1.149 | 0.6658 | 1.072 |
| Quadruplet-TSST | 5.3 | 8.6 | 802 | 53.9 | 1.116 | 0.6510 | 1.057 |
| Quadruplet-STTS | 5.3 | 8.6 | 858 | 54.9 | 1.118 | 0.6507 | 1.057 |
WeatherBench. Climate prediction is a critical challenge in spatiotemporal predictive learning. The WeatherBench (Rasp et al., 2020) dataset provides a comprehensive global weather forecasting resource, covering various climatic factors. In our experiments, we utilize WeatherBench-S, a single-variable setup where each climatic factor is trained independently. We focus on temperature prediction at a 5.625∘ resolution ( grid points). The model is trained on data spanning 2010-2015, validated on data from 2016, and tested on data from 2017-2018, all with a one-hour temporal interval. We input the first 12 frames and predict the subsequent 12 frames in this setting.
@@ 4.5 Ablation Study and Discussion
We conduct ablation studies on PredFormer model design and summarize the results in Tab 7 and Tab 8. We choose the best Triplet-STS model on TaxiBJ, and the best Fac-T-S model on WeatherBench as baselines.
@@
Model Regularization. Pure transformer architectures like ViT generally require large datasets for effective training, and overfitting can become challenging when applied to smaller datasets. In our experiments, overfitting is noticeable on WeatherBench and TaxiBJ. We experiment with different regularization techniques in Tab 8 and find that both dropout(DP) and stochastic depth (SD) individually improve performance compared to no regularization. However, the combination of the two provides the best results. Unlike conventional ViT practices, which use a linearly scaled drop path rate across different depths, a uniform drop path rate performs significantly better for our tasks.
@@
Discussion for PredFormer Variants. Despite our in-depth analysis of the spatiotemporal decomposition, the optimal model is not definite due to the different spatiotemporal dependent properties of the datasets.
We recommend starting with the Quadruplet-TSST model for diverse video prediction tasks, which consistently performs well across datasets and configurations. Use M Quadruplet-TSST layers and experiment with models having a total of 4M GTBs to identify the optimal configuration. Then, explore Triplet-TST and Triplet-STS with M layers to find spatial and temporal dependencies. Unlike the SimVP framework, which adjusts hidden dimensions and block numbers separately for spatial encoder-decoder and temporal translator, PredFormer uses fixed hyperparameters for spatial and temporal GTBs, leveraging the scalability of the Transformer architecture. By simply adjusting the number of PredFormer layers, optimal results can be achieved with minimal tuning. We provide further theoretical analysis in the Appendix.
@@ 5 Conclusion
In this paper, we introduce PredFormer, a pure Transformer-based framework for video prediction, and systematically study how different spatio-temporal factorization patterns affect efficiency and accuracy. Across four standard benchmarks, PredFormer achieves strong performance while maintaining favorable FLOPs, FPS, and memory usage compared to prior CNN- and RNN-based models, providing a competitive recurrent-free and convolution-free alternative. On the four benchmarks we study, our experiments suggest several dataset-dependent tendencies: interleaved spatio-temporal architectures generally offer the best balance between cost and accuracy; temporal-first factorizations often work particularly well for long-horizon forecasting, while other interleaved patterns can be preferable when the spatial structure is richer or the horizon is shorter; combining dropout with uniform stochastic depth is especially effective on overfitting-prone datasets; and absolute positional encoding is consistently more robust than learnable alternatives.
Appendix A Appendix
A.1 Problem Definition